A Bibliography of Publications in Notices of the American Mathematical Society (1960–1969)

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

15 August 2024 Version 1.00

Title word cross-reference

2^{2m+1} [Knu64b, Knu62]. n [Knu64a]. V = L [Jen68]. -th [Knu64a]. Adaptive [Hol62]. Addition [Knu64a]. Automata [Bur70, vNB66]. Bose [KF67]. Cellular [Bur70]. chains [Knu64a]. combinatorial [HU60]. constructions [KF67]. Constructive [BK69]. Desarguesian [Knu62, Knu64b].

Dimensions [SU60]. enumeration [BK69]. Essays [Bur70]. evaluation [Knu64a]. Figures [SU60]. Growth [SU60, HU60]. Hilbert [Tau69]. hypothesis [Jen68]. II [Tau69]. Improved [KF67]. incompatible [Jen68]. Logical [Hol62]. Nelson [KF67]. Non [Knu62, Knu64b]. Non-Desarguesian [Knu62, Knu64b]. order [Knu62, Knu64b]. Outline [Hol62].

partitions [BK69]. Patterns [SU60, HU60]. plane [BK69]. planes [Knu62, Knu64b]. powers [Knu64a]. problem [KF67]. problems [HU60].

Reproducing [vNB66].

Self [vNB66]. Self-Reproducing [vNB66]. some [HU60]. sorting [KF67]. Souslin [Jen68]. Systems [Hol62].

th [Knu64a]. Theorem [Tau69]. Theory [Hol62, vNB66]. Two [SU60].

References

Bender:1969:CEP

[BK69] Edward A. Bender and Donald E.

Knuth. Constructive enumeration of plane partitions. Notices of the American Mathematical Society, 16 (4):659, June 1969. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Burks:1970:ECA

[Bur70] Arthur W. Burks. Essays on Cellular Automata. University of Illinois Press, Urbana, IL, USA, 1970. xxvi + 375 pp. LCCN QA267.5.S4 B87. A classic collection of papers on cellular automata intended as a companion to von Neumann’s Theory of Self Reproducing Automata [vNB66]. Includes papers by Burks, Thatcher, Moore, Myhill, Ulam, and Holland. See [Hol62].

Holland:1962:OLT

[Hol62] John H. Holland. Outline for a logical theory of adaptive systems.

Journal of the ACM, 9(3):297–314, July 1962. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). Early work on genetic algorithms. Reprinted in [Bur70].

Holladay:1960:SCP

[HU60] J. C. Holladay and Stanisław M. Ulam. On some combinatorial problems in patterns of growth, I. Notices of the American Mathematical Society, 7(2):234, April 1960. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Jensen:1968:SHI

[Jen68] R. B. Jensen. Souslin’s hypothesis is incompatible with V = L. Notices of the American Mathematical Society, 15(6):935, October 1968. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Knuth:1967:ICB

[KF67] Donald E. Knuth and R. W. Floyd. Improved constructions for the Bose–Nelson sorting problem. Notices of the American Mathematical Society, 14(2):283, February 1967. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Knuth:1962:NDP

[Knu62] Donald E. Knuth. Non-Desarguesian planes of order 2^{2m+1}. Notices of the American Mathematical Society, 9(3):218, June 1962. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Knuth:1964:ACE

[Knu64a] Donald E. Knuth. Addition chains and the evaluation of n-th powers. Notices of the American Mathematical Society, 11(2):230–231, February 1964. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Knuth:1964:NDP

[Knu64b] Donald E. Knuth. Non-Desarguesian planes of order 2^{2m+1}. Notices of the American Mathematical Society, 11 (4):445–446, June 1964. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Schrandt:1960:PGF

[SU60] R. G. Schrandt and S. M. Ulam.

On patterns of growth of figures in two dimensions. Notices of the American Mathematical Society, 7 (5):642, October 1960. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). Early studies of pattern development in cellular automata, including competitions between two kinds of patterns. Reprinted in [Bur70].

Taussky:1969:HTI

[Tau69] Olga Taussky. Hilbert’s theorem 94.

II. Notices of the American Mathematical Society, 16(2):407, February 1969. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

vonNeumann:1966:TSR

[vNB66] John von Neumann and Arthur W. (Arthur Walter) Burks, editors. Theory of Self-Reproducing Automata. University of Illinois Press, Urbana, IL, USA, 1966. xix + 388 pp. LCCN QA267 .V55. Von Neumann’s work on selfreproducing automata, completed and edited after his death by Arthur Burks. Also includes transcripts of von Neumann’s 1949 University of Illinois lectures on the “Theory and Organization of Complicated Automata”. See [Bur70].